# Egor Cherepanov1,2Nikita Kachaev1Alexey K. Kovalev1,2Aleksandr I. Panov1,2 1AIRI, Moscow, Russia  2MIPT, Dolgoprudny, Russia

### v1

@@ 4 Classification of memory-intensive tasks
To address these challenges, we propose a systematic approach to memory evaluation in RL. Given the impracticality of testing agents on every possible memory-intensive environment, we aim to identify a minimal diagnostic set that comprehensively covers different memory requirements. Drawing from established research in developmental psychology and cognitive science, where similar memory challenges have been extensively studied in humans, we develop a categorization framework consisting of four distinct memory task classes, detailed in Subsection 4.2.
@@ 4.1 Memory: From Cognitive Science to RL
In developmental psychology and cognitive science, memory is classified into categories based on cognitive processes. Key concepts include object permanence (Piaget, 1952), which involves remembering the existence of objects out of sight, and categorical perception (Liberman et al., 1957), where objects are grouped based on attributes like color or shape. Working memory (Baddeley, 1992) and memory span (Daneman & Carpenter, 1980) refer to the ability to hold and manipulate information over time, while causal reasoning (Kuhn, 2012) and transitive inference (Heckers et al., 2004) involve understanding cause-and-effect relationships and deducing hidden relationships, respectively.
The RL field has attempted to utilize these concepts in the design of specific memory-intensive environments Fortunato et al. (2020); Lampinen et al. (2021), but these have been limited at the task design level. Of particular interest, however, is how existing memory-intensive tasks can be categorized using these concepts to develop a benchmark on which to test the greatest number of memory capabilities of memory-enhanced agents, and it is this problem that we address in this paper. Thus, we aim to provide a balanced framework that covers important aspects of memory for real-world applications while maintaining practical simplicity (see  Figure 3).
@@ 4.2 Taxonomy of Memory Tasks
    Object Memory. Tasks that evaluate an agent’s ability to maintain object-related information over time, particularly when objects become temporarily unobservable. These tasks align with the cognitive concept of object permanence, requiring agents to track object properties when occluded, maintain object state representations, and recognize encountered objects.
@@
    Spatial Memory. Tasks focused on environmental awareness and navigation, where agents must remember object locations, maintain mental maps of environment layouts, and navigate based on previously observed spatial information.
@@
    Sequential Memory. Tasks that test an agent’s ability to process and utilize temporally ordered information, similar to human serial recall and working memory. These tasks require remembering action sequences, maintaining order-dependent information, and using past decisions to inform future actions.
@@
    Memory Capacity. Tasks that challenge an agent’s ability to manage multiple pieces of information simultaneously, analogous to human memory span. These tasks evaluate information retention limits and multi-task information processing.
@@ 5 MIKASA-Base
Table 3: Analysis of established robotics frameworks with manipulation tasks, comparing their support for memory-intensive tasks. $\dagger$ – excluding Franka Kitchen.
Robotics Framework with Manipulation Tasks Memory Tasks Manipulation Atomic Low-level actions MIKASA-Robo (Ours) ManiSkill3 (Tao et al., 2024) ManiSkill-HAB (Shukla et al., 2024) RoboCasa (Nasiriany et al., 2024) Gymnasium-Robotics (de Lazcano et al., 2024) BEHAVIOR-1K (Li et al., 2024) ARNOLD (Gong et al., 2023) iGibson 2.0 (Li et al., 2022) VIMA (Jiang et al., 2022) Isaac Sim (Makoviychuk et al., 2021) panda-gym (Gallouédec et al., 2021) Habitat 2.0 (Szot et al., 2021) Meta-World (Yu et al., 2020) CausalWorld (Ahmed et al., 2020) RLBench (James et al., 2020) robosuite (Zhu et al., 2020b) dm_control (Tunyasuvunakool et al., 2020) Franka Kitchen (Gupta et al., 2019) SURREAL (Fan et al., 2018) AI2-THOR (Kolve et al., 2017)
@@
The RL domain currently lacks standardized benchmarks for evaluating agents’ memory capabilities. While numerous memory-intensive environments exist, their dispersion across different research projects makes systematic comparison challenging. Moreover, existing frameworks often focus on narrow aspects of memory, failing to capture the diverse memory requirements found in real-world applications. To address these limitations, we introduce MIKASA-Base, a unified benchmark that systematically evaluates memory capabilities across diverse tasks while maintaining practical simplicity.
@@
Our benchmark follows key design principles that ensure comprehensive evaluation of memory capabilities. To isolate memory mechanisms from other learning challenges, MIKASA-Base implements a two-tiered task structure. The first tier consists of diagnostic vector-based environments, enabling direct validation of specific memory mechanisms in atomic tasks. The second tier comprises complex image-based environments that introduce additional challenges through 2D observation processing, more closely approximating real-world scenarios. This hierarchical approach allows researchers to first validate fundamental memory capabilities before progressing to more sophisticated tasks.
@@
Building upon our taxonomy presented in Subsection 4.2, we conducted a systematic analysis of existing open-source memory-intensive environments from Table 2. Our analysis revealed four distinct classes of memory tasks. This classification enabled us to identify a minimal yet representative set of environments that spans the large spectrum of memory utilization patterns, from object permanence to sequential decision-making, while maintaining practical simplicity. Detailed descriptions of all considered environments are provided in Appendix F, while Table 5 in the Appendix presents an analysis of MIKASA-Base memory-intensive environments.
We unified these environments under the Gymnasium API (Towers et al., 2024), enabling seamless integration with existing RL tools (see Table 4). This standardization facilitates direct architectural comparisons. Implementation details are provided in Appendix B.
Table 4: Recommended memory-intensive environments for comprehensive agent evaluation.
Memory Type Diagnostic Tasks Complex Tasks Object Memory Passive T-Maze (Ni et al., 2023) ViZDoom-Two-Colors (Sorokin et al., 2022) Spatial Memory POPGym Labyrinth (Morad et al., 2023) Memory Maze (Pasukonis et al., 2022) Sequential Memory POPGym Autoencode (Morad et al., 2023) Ballet (Lampinen et al., 2021) Memory Capacity Memory Cards (Esslinger et al., 2022) MemoryGym Mortar Mayhem (Pleines et al., 2023)
To evaluate agents in realistic memory-intensive scenarios, we introduce our MIKASA-Robo benchmark (Section 6). This benchmark provides a suite of robotic manipulation tasks that systematically assess all four memory types in practical, real-world-inspired contexts.
MIKASA-Base standardizes memory evaluation in RL through organized environment selection and structured task progression. Through its carefully curated environment selection and hierarchical structure, MIKASA-Base enables systematic evaluation of memory-enhanced architectures, facilitates direct comparison between different memory mechanisms, and provides a clear progression path from fundamental to complex memory tasks. This structured approach allows precise identification of memory-related limitations in RL agents while maintaining practical utility.
@@ 6 MIKASA-Robo
The landscape of robotic manipulation frameworks reveals significant limitations in addressing memory-intensive tasks. First, while partial observability is extensively studied in navigation tasks, manipulation scenarios are predominantly evaluated under full observability, with memory requirements receiving limited attention (see Table 3). Second, among the few frameworks that incorporate memory-intensive manipulation tasks, significant limitations exist. BEHAVIOR-1k (Li et al., 2024) and iGibson 2.0 (Li et al., 2022) employ highly complex, non-atomic tasks that obscure the evaluation of specific memory mechanisms. Similarly, VIMA (Jiang et al., 2022) relies on high-level actions that inadequately capture memory performance over extended time horizons. To the best of our knowledge, there are no benchmarks specifically designed to evaluate memory in RL in the robotic manipulation domain. To fill this gap, we introduce the MIKASA-Robo framework for the RL.
6.1 MIKASA-Robo Benchmark
@@
Building upon the robust foundation of ManiSkill3 framework (Tao et al., 2024), our benchmark leverages its efficient parallel GPU-based training capabilities to create and evaluate these tasks.
@@ 6.2 MIKASA-Robo Manifestation
In designing the tasks, we drew inspiration from the four memory types identified in our classification framework (Subsection 4.2). We developed 32 tasks across 12 categories of robotic tabletop manipulation, each targeting specific aspects of object memory, spatial memory, sequential memory, and memory capacity. These tasks feature varying levels of complexity, allowing for systematic evaluation of different memory mechanisms. For instance, some tasks test object permanence by requiring the agent to track occluded objects, while others challenge sequential memory by requiring the reproduction of a strict order of actions. A summary of these tasks and their corresponding memory types is provided in Table 1, with detailed descriptions in Appendix E.
@@
Figure 5: PPO with MLP and LSTM backbones trained in RGB+joints mode on the RememberColor-v0 environment with dense rewards. Both architectures fail to solve medium and high complexity tasks.
@@
6.3 Performance of classic baselines on MIKASA-Robo benchmark
@@
For our experimental evaluation, we selected PPO (Schulman et al., 2017) with two backbone architectures: Multilayer Perceptron (MLP) and Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997). The MLP variant serves as a memory-less baseline, while LSTM represents a widely-adopted memory mechanism in RL, known for its effectiveness in solving POMDPs (Ni et al., 2021). This choice of architectures enables direct comparison between memory-less and memory-enhanced agents while validating our benchmark’s ability to assess memory. We focus specifically on these fundamental architectures as they align with our primary goal of benchmark validation rather than comprehensive algorithm comparison. To demonstrate that all proposed environments are solvable with 100% success rate (SR), we trained a PPO-MLP agent using state mode, where it had full access to system information. Results for the demo environments are presented in Figure 5, with additional results for all tasks in Appendix D.
@@
Training under the RGB+joints mode with dense rewards reveals the memory-intensive nature of our tasks. Using the RememberColor-v0 task as an example, PPO-LSTM demonstrates superior performance compared to PPO-MLP when distinguishing between three colors (see Figure 5). However, both agents’ success rates drop dramatically to near-zero as the task complexity increases to five or nine colors. Moreover, under sparse reward conditions, both architectures fail to solve even the three-color variant (see Appendix, Figure 9). These results validate our benchmark’s effectiveness in evaluating agents’ memory, showing clear performance degradation as memory demands increase.
Our baseline experiments reveal key insights: (1) the proposed tasks are inherently solvable, as demonstrated by perfect performance in state mode; (2) the tasks effectively challenge memory capabilities, shown by the performance gap between memory-less (MLP) and memory-enhanced (LSTM) architectures; and (3) primitive memory mechanisms show limitations as task complexity increases, particularly under sparse rewards. These findings validate MIKASA-Robo as an effective benchmark for evaluating and developing memory-enhanced RL agents in robotic tasks.
@@ 7 Conclusion
In this work, we addressed several critical gaps in memory-enhanced RL research: the lack of standardized agents’ evaluation methods, the absence of a unified taxonomy for memory tasks, and the disconnect between abstract memory challenges and practical robotics applications. We addressed these through three key contributions. First, we developed a comprehensive classification framework that categorizes memory tasks into four distinct classes: object memory, spatial memory, sequential memory, and memory capacity. This taxonomy provides a structured approach to understanding and evaluating different aspects of memory in RL agents. Second, we introduced MIKASA-Base, a unified benchmark that consolidates diverse memory-intensive environments into a single, standardized framework. By carefully selecting representative tasks from each memory category, our benchmark enables systematic comparison and evaluation of memory-enhanced RL agents across a broad spectrum of memory challenges. Third, we presented MIKASA-Robo, a novel benchmark comprising 32 carefully designed memory-intensive tasks for robotic manipulation, which bridges the gap between abstract memory challenges and practical robotics applications. We hope that our contributions will serve as a foundation for future research in memory-enhanced RL, accelerating the development of more capable and reliable autonomous systems for real-world applications.
Acknowledgments.
The research was carried out with the support of the Sber Robotics Center as part of the work at the MIPT Research Center for Applied Artificial Intelligence Systems.

### v2

@@ 4 Classification of memory-intensive tasks
To address these challenges, we propose a systematic approach to memory evaluation in RL. Drawing from established research in developmental psychology and cognitive science, where similar memory challenges have been extensively studied in humans, we develop a categorization framework consisting of four distinct memory task classes, detailed in Subsection 4.2.
@@ 4.1 Memory: From Cognitive Science to RL
In developmental psychology and cognitive science, memory is classified into categories based on cognitive processes. Key concepts include object permanence [74], which involves remembering the existence of objects out of sight, and categorical perception [60], where objects are grouped based on attributes like color or shape. Working memory [4] and memory span [16] refer to the ability to hold and manipulate information over time, while causal reasoning [50] and transitive inference [35] involve understanding cause-and-effect relationships and deducing hidden relationships, respectively.
The RL field has attempted to utilize these concepts in the design of specific memory-intensive environments [22, 54], but these have been limited at the task design level. Of particular interest, however, is how existing memory-intensive tasks can be categorized using these concepts to develop a benchmark on which to test the greatest number of memory capabilities of memory-enhanced agents, and it is this problem that we address in this paper. Thus, we aim to provide a balanced framework that covers important aspects of memory for real-world applications while maintaining practical simplicity (see  Figure 3).
@@ 4.2 Taxonomy of Memory Tasks
    Object Memory. Tasks that evaluate an agent’s ability to maintain object-related information over time, particularly when objects become temporarily unobservable. These tasks align with the cognitive concept of object permanence, requiring agents to track object properties when occluded, maintain object state representations, and recognize encountered objects. Example: a robot remembers which fruit it put in the fridge.
@@
    Spatial Memory. Tasks focused on environmental awareness and navigation, where agents must remember object locations, maintain mental maps of environment layouts, and navigate based on previously observed spatial information. Example: the robot remembers the position of a mug it moved while cleaning and returns it to its place.
@@
    Sequential Memory. Tasks that test an agent’s ability to process and utilize temporally ordered information, similar to human serial recall and working memory. These tasks require remembering action sequences, maintaining order-dependent information, and using past decisions to inform future actions. Example: a robot memorizes the order of the ingredients it has added to a soup.
@@
    Memory Capacity. Tasks that challenge an agent’s ability to manage multiple pieces of information simultaneously, analogous to human memory span. These tasks evaluate information retention limits and multi-task information processing. Example: a robot is able to memorize the positions of several different objects while cleaning a table.
@@ 5 MIKASA-Base
Table 3: Analysis of established robotics frameworks with manipulation tasks, comparing their support for memory-intensive tasks. $\dagger$ – excluding Franka Kitchen. * – concurrent work with three memory tasks with only one type of memory.
Robotics Framework with Manipulation Tasks Memory Tasks Manipulation Atomic Low-level actions MIKASA-Robo (Ours) MemoryBench [21] ManiSkill3 [87] ManiSkill-HAB [81] FetchBench [32] RoboCasa [66] Gymnasium-Robotics [17] BEHAVIOR-1K [59] ARNOLD [24] iGibson 2.0 [58] VIMA [41] Isaac Sim [63] panda-gym [23] Habitat 2.0 [86] Meta-World [96] CausalWorld [1] RLBench [40] robosuite [102] dm_control [91] Franka Kitchen [29] SURREAL [20] AI2-THOR [49]
@@
Despite the importance of memory in decision-making, the RL community lacks standardized tools for benchmarking memory capabilities. Existing studies typically introduce bespoke environments tailored to their proposed algorithms, leading to fragmentation and limited comparability across works (see Table 2). Moreover, many popular memory benchmarks focus narrowly on specific memory types, overlooking the diversity of memory demands found in real-world applications. To address this gap, we introduce MIKASA-Base, a unified benchmark that consolidates widely used open-source memory-intensive environments under a common Gym-like API. Our goal is to streamline reproducibility, support fair comparisons, and promote systematic evaluation of memory in RL.
@@
MIKASA-Base is designed around core principles that support rigorous and interpretable evaluation of memory in RL. To disentangle memory from unrelated challenges, we organize tasks into two tiers. The first tier consists of diagnostic vector-based environments that isolate specific memory mechanisms. The second tier includes complex image-based tasks that incorporate realistic perception challenges, thus more closely resembling real-world settings. This hierarchical structure enables researchers to validate memory capabilities incrementally – from atomic reasoning to high-dimensional sensory input.
@@
Building on our taxonomy from Subsection 4.2, we systematically reviewed open-source memory benchmarks and categorized their tasks into four distinct types of memory usage. We selected a diverse yet representative subset of environments to cover this taxonomy – ranging from object permanence to sequential planning. All selected tasks are unified under a single, consistent API. Descriptions are provided in Appendix I, and an overview of MIKASA-Base tasks appears in Table 7. This consolidation supports architectural ablations, direct comparison of methods, and simplified evaluation pipelines. Implementation details can be found in Appendix C.
MIKASA-Base provides the first systematic and unified benchmark for evaluating memory in RL. It mitigates fragmentation by standardizing task access and evaluation, and its structured progression enables precise attribution of memory-related agent failures. By covering a broad spectrum of memory challenges within a common framework, MIKASA-Base offers a foundation for robust, reproducible research in memory-centric RL.
@@ 6 MIKASA-Robo
The landscape of robotic manipulation frameworks reveals significant limitations in addressing memory-intensive tasks. While partial observability is well-studied in navigation, manipulation scenarios are still predominantly evaluated under full observability, with limited focus on memory demands (see Table 3). Among frameworks that do consider memory, BEHAVIOR-1k [59] and iGibson 2.0 [58] include highly complex, non-atomic tasks, which obscure the evaluation of specific memory mechanisms. VIMA [41] relies on high-level action abstractions, limiting temporal memory assessment. To address these gaps, we introduce MIKASA-Robo, a benchmark specifically designed to evaluate diverse memory skills in robotic manipulation through well-isolated, fine-grained tasks.
Concurrently with our work, Fang et al. [21] proposed MemoryBench, a benchmark focused on spatial memory with three robotic tasks. In contrast, MIKASA-Robo spans four memory categories and 32 tasks, enabling broader and more systematic evaluation of memory mechanisms in RL agents.
@@
Building upon the robust foundation of ManiSkill3 framework [87], our benchmark leverages its efficient parallel GPU-based training capabilities to create and evaluate these tasks.
@@ 6.1 MIKASA-Robo Manifestation
In designing the tasks, we drew inspiration from the four memory types identified in our classification framework (Subsection 4.2). We developed 32 tasks across 12 categories of robotic tabletop manipulation, each targeting specific aspects of object memory, spatial memory, sequential memory, and memory capacity. These tasks feature varying levels of complexity, allowing for systematic evaluation of different memory mechanisms. For instance, some tasks test object permanence by requiring the agent to track occluded objects, while others challenge sequential memory by requiring the reproduction of a strict order of actions. A summary of these tasks and their corresponding memory types is provided in Table 1, with detailed descriptions in Appendix H.
@@
Figure 5: Online RL baselines with MLP and LSTM backbones trained in RGB+joints mode on the RememberColor-v0 environment with dense rewards. Both architectures fail to solve medium and high complexity tasks.
@@
Figure 6: Results of Offline RL baselines with memory (RATE, DT) and without memory (BC-MLP, CQL-MLP, DP) on all 32 MIKASA-Robo tasks. Training was performed in RGB mode with sparse rewards (success condition).
6.2 Online RL baselines
For the experimental evaluation, we chose on-policy Proximal Policy Optimization (PPO, [79]) with two underlying architectures: Multilayer Perceptron (MLP) and Long Short-Term Memory (LSTM, [37]), as well as popular in robotics off-policy Soft Actor-Critic (SAC, [30]) and model-based Temporal Difference Learning for Model Predictive Control (TD-MPC2, [33]).
@@ 6.2 Online RL baselines
The MLP variant serves as a memory-less baseline, while LSTM represents a widely-adopted memory mechanism in RL, known for its effectiveness in solving POMDPs [67]. This choice of architectures enables direct comparison between memory-less and memory-enhanced agents while validating our benchmark’s ability to assess memory. We focus specifically on these fundamental architectures as they align with our primary goal of benchmark validation rather than comprehensive algorithm comparison. To demonstrate that all proposed environments are solvable with 100% success rate (SR), we trained a PPO-MLP agent using state mode, where it had full access to system information. Results for select tasks are shown in Figure 5; full results are in Appendix F.
@@
Training under the RGB+joints mode with dense rewards reveals the memory-intensive nature of our tasks. Using the RememberColor-v0 task as an example, PPO-LSTM demonstrates superior performance compared to PPO-MLP when distinguishing between three colors (see Figure 5). However, both agents’ success rates drop dramatically to near-zero as the task complexity increases to five or nine colors. Moreover, under sparse reward conditions, both architectures fail to solve even the three-color variant (see Appendix FFigure 10). Additionally, our findings indicate that, while SAC and TD-MPC2 exhibit higher sample efficiency compared to PPO-MLP, when faced with more complex challenges, the lack of an explicit memory mechanism becomes a critical shortcoming, resulting in low performance, which also emphasizes the inappropriateness of algorithms common in the robotics community for memory-intensive tasks. These results validate our benchmark’s effectiveness in evaluating agents’ memory, showing clear performance degradation as memory demands increase.
6.3 Offline RL baselines
Since dense rewards are typically not available in the real world, it is of particular interest to train on sparse rewards represented as a binary flag of a successfully completed episode. Whereas models with online learning are extremely hard to handle in this setting, we also conducted experiments with five Offline RL models: Decision Transformer (DT) [8]) and Recurrent Action Transformer with Memory (RATE) [10]) based on the Transformer architecture, Standard Behavioral Cloning (BC) and Conservative Q-Learning (CQL) [51]) with MLP backbones, as well as Diffusion Policy (DP) [13]) – a recent and popular approach in robotic manipulation that leverages diffusion models for direct action prediction.
Experimental results with Offline RL models trained using two RGB camera views and sparse rewards are presented in Figure 6. As can be seen from Figure 6, none of the models – including those explicitly designed for sequence modeling – were able to successfully solve the majority of MIKASA-Robo tasks, demonstrating the challenge posed by the benchmark. Training was conducted using datasets consisting of 1000 successful trajectories per task (see Appendix B for details).
Notably, none of the evaluated models were able to solve tasks requiring high Memory Capacity or Sequential Memory, further underscoring their complexity. More detailed results for Offline RL algorithms are presented in Appendix, Table 6.
6.4 VLA baselines
To investigate the capabilities of state-of-the-art Visual-Language-Action (VLA) models in memory-intensive robotic tasks, we selected two representative baselines: Octo [88] and OpenVLA [47]. Although neither model explicitly claims to implement sophisticated memory mechanisms, these experiments provide valuable insights into the current state of memory capabilities in VLA agents.
Octo is a transformer with diffusion heads trained from scratch on 25 Open X-Embodiment datasets [15]; in our experiments, only the readout heads were fine-tuned using the full pretrained context length of 10 and action chunk size $K=4$. OpenVLA uses a Prismatic-7B backbone [46], fine-tuned for action prediction with LoRA adapters, action chunking, and an $L_{1}$ loss [48]. We tested chunk sizes $K=4$ and $K=8$. Both models were trained on 250 expert trajectories per task, using $128\times 128$ RGB image pairs (base and wrist views) and end-effector control (see Appendix D).
Table 4: Performance of VLA models on selected memory-intensive tasks from the MIKASA-Robo benchmark. Reported values denote average success rates over 100 evaluation episodes (mean $\pm$ sem). Tasks include spatial reasoning (ShellGameTouch, InterceptMedium) and color-based memory retrieval (RememberColor3/5/9).
Model ShellGameTouch InterceptMedium RememberColor3 RememberColor5 RememberColor9 Octo-small 0.46 $\pm$ 0.05 0.39 $\pm$ 0.04 0.45 $\pm$ 0.06 0.17 $\pm$ 0.03 0.11 $\pm$ 0.03 OpenVLA ($K{=}4$) 0.12 $\pm$ 0.05 0.06 $\pm$ 0.02 0.21 $\pm$ 0.00 0.09 $\pm$ 0.02 0.08 $\pm$ 0.02 OpenVLA ($K{=}8$) 0.47 $\pm$ 0.05 0.14 $\pm$ 0.03 0.59 $\pm$ 0.04 0.16 $\pm$ 0.03 0.06 $\pm$ 0.02
Experimental results (Table 4) reveal notable trends. Octo (context size 10) outperforms random on simpler tasks, suggesting some innate memory capacity, but its performance degrades with task complexity, indicating limited scalability. OpenVLA shows contrasting behavior across chunk sizes: with $K=8$, it exceeds random on tasks like RememberColor3 and ShellGameTouch, despite lacking step-wise history. However, performance drops sharply on harder tasks. With $K=4$, it fails across the board. These results suggest that larger chunk sizes can help bypass explicit memory by generating full trajectories from early cues, but this strategy fails with smaller chunks, where initial correct actions often give way to confusion. Thus, action chunking offers limited compensation for the absence of a true memory mechanism.
The sharp decline in performance on higher-complexity tasks underscores the necessity for dedicated memory architectures and validates the importance of the multi-difficulty hierarchy in MIKASA-Robo to prevent such “shortcuts”. Our experiments with Octo and OpenVLA highlights a critical gap in current VLA models: the absence of effective long-term memory leads to brittle performance on tasks demanding strong memory capabilities. These experiments not only illuminate present limitations but also reinforce the value of the MIKASA-Robo benchmark.
7 Limitations
While our benchmark provides a comprehensive evaluation framework, some limitations remain. In particular, the performance of Octo and OpenVLA may not reflect their full potential, as we performed limited fine-tuning due to computational constraints. Future work could explore more extensive adaptation of large VLA models within MIKASA to better assess their memory capabilities. Additionally, while MIKASA covers a broad range of memory challenges, further extensions could incorporate tasks with longer temporal dependencies or meta-RL.
@@ 8 Conclusion
We present MIKASA, a unified benchmark suite for evaluating memory in RL. Our work addresses key gaps in the field by introducing: (1) a taxonomy of memory types – object, spatial, sequential, and capacity; (2) MIKASA-Base, a standardized collection of open-source memory tasks; (3) MIKASA-Robo, a suite of 32 robotic manipulation tasks targeting diverse memory demands; and (4) accompanying offline datasets to support reproducible evaluation. Experiments with online, offline, and VLA agents reveal that current methods struggle with many tasks, highlighting the need for better memory architectures. MIKASA aims to guide and accelerate progress in memory-intensive RL for real-world applications. The MIKASA-Robo suite is publicly available and can be easily installed via pip install mikasa-robo-suite.